Perform the RF communication simulation for FSK modulation, for the modulation and transmission of the information signal with the carrier on the proteus, and for the re-acquisition with a coherent type structure on the receiving side. Display communication signals. (You can use the carrier and information signal at the frequency you want)
Answer #1

1. 3 bits

In 8PSK modulator, every 3 input data bits are mapped to one complex-valued 8PSK symbol.

2. Spectrum is about 0 Hz

A baseband signal is one which has a spectrum from 0 Hz to some fc Hz where fc is the cut off frequency. It is a low pass signal whose frequency is close to zero with respect to the highest frequency of the same signal.

3. It needs a smaller bandwidth.

DPSK stands for Differential Phase Shift Keying. It is the version of BPSK. In DPSK, there is no absolute carrier phase reference, instead of transmitted signal itself used as the phase reference.

The advantage of DPSK receiver is it’s simpler to implement because it does not require a coherent carrier recovery subsystem also carrier recovery circuitry is not required.

4. 5 kHz

The bit period in the baseband signal is T= 1/ 10000 seconds

The bandwidth of the baseband signal= 1/2T

B= 2×1000


answered by: ANURANJAN SARSAM